OpenStudy (anonymous):

hamer el purchased a leather easy chair and ottoman, black steel end table, and a black contemporary floor lamp at a total of $308.90 with a 6.0% sales tax. what is the total purchase price?

OpenStudy (anonymous):

\[1.06\times P = 308.90\] \[P=308.90\div 1.06\]

OpenStudy (anonymous):

i get about $291.5\] rounded

OpenStudy (anonymous):

unless it means what was the price AFTER tax in which case it is just \[308.90\times 1.06\]

OpenStudy (anonymous):

total purchase price will be = 308.90 + 0.06(308.90) = 308.90 + 18.53 = 327.43

OpenStudy (anonymous):

thank you again ha.

OpenStudy (anonymous):

welcome cody o( :{)